(4 sqrt 5 - 2 sqrt 3) (4 sqrt 5 + 2 sqrt 3) thank you for all your help!
Answer by checkley71(8403)

(4SQRT5-2SQRT3)(4SQRT5+2SQRT3)
4SQRT5*4SQRT5=16*5=90
4SQRT5*2SQRT3=8*SQRT15
-2SQRT3*4SQRT5=-8SQRT15
-2SQRT3*2SQRT3=-4*3
NOW WE HAVE 90+8SQRT15-8SQRT15-12 SQRT15 CANCELLS THUS WE HAVE
90-12=78